A household cleaner contains 4.25 g of ammonia in 0.50 L of solution. There are 17.01 g in 1 mol of ammonia. What is the concentration of ammonia in mol/L?


1
Expert's answer
2021-06-30T03:57:42-0400

Solution

no of moles"=\frac{mass}{molar mass}=\frac{4.25}{17.01}"


= 0.2499 moles


molarity"=\frac{no of moles}{volume inL}= \frac{0.2499}{0.5}"


=0.4998 mol/L